OpenStudy (anonymous):

Pamela is 13 years older than Jiri. The sum of their ages is 73 . What is Jiri's age?

OpenStudy (anonymous):

73-13=60 60/2=30 30 is jiri age

OpenStudy (anonymous):

43 is pamela

OpenStudy (anonymous):

(13+J)+J=73 2J = 73 -13 J = 60/2 J=30 Jiri = 30 Pamela is = 30+13=43

OpenStudy (anonymous):

let the age of pamela = P and the age of jiri = J P = 13 + J P + J = 73 then substitute the value of P ( 13 + j )+J = 73 13 + 2J = 73 2J = 60 J = 30 thus jiri's age is 30